# LDMol: Text-Conditioned Molecule Diffusion Model Leveraging Chemically Informative Latent Space

### v1

@@ 3.2 Text-conditioned latent diffusion model
As shown in Figure 2-(c), our diffusion model learns the conditional distribution of the SMILES latent $z$ whose dimension size is $\{L\times d_{z}\}$. In the training phase, a molecule in the training data is mapped to the latent $z$ and applied a forward noising process into $z_{t}$ with randomly sampled diffusion timestep $t$ and injected noise $\epsilon$. A diffusion model predicts the injected noise from $z_{t}$, conditioned by the paired text description via a frozen external text encoder. In the inference phase, the diffusion model iteratively generates a new latent sample $z^{\prime}$ from a given text condition, which is then decoded to a molecule via the SMILES decoder.
Since most contributions to diffusion models were made in the image domain, most off-the-shell diffusion models have the architecture of convolution-based Unet [40]. However, to step beyond the spatial inductive bias and make the model applicable for more general data modality, diffusion models using transformer architectures are proposed. One of the most successful approaches among them is DiT [35], which is the model we’ve employed. Specifically, we utilized a DiTbase model with two major modifications: first, DiT’s two-dimensional attributes (e.g., patch embedding) were modified accordingly to handle one-dimensional latent. Second, cross-attention layers for external conditions were added after the self-attention layers to receive high-dimensional text conditions.
Text-based molecule generation requires a text encoder to process natural language conditions. Existing text-based molecule generation models trained their text encoder from scratch [36], or utilize a separate encoder model pre-trained on scientific domain corpora [3]. In this work, we took the encoder part of MolT5 [10] as our text encoder.
@@ 3.3 Implementation details
The pre-training of the SMILES encoder and the corresponding decoder was done with 10,000,000 general molecules from PubChem [24]. To ensure enough batch size for negative samples [15], we build a memory queue that stores $Q$ recent input and use them for the encoder training. To be fed into the autoencoder, an input SMILES must be tokenized as a sequence of tokens in a pre-defined vocabulary. The tokenizer vocabulary consists of 300 tokens, which were obtained from the pre-training data SMILES corpus using the BPE algorithm [11]. We only used SMILES data that does not exceed a fixed maximum token length $L$.
To train the text-conditioned latent diffusion model, we gathered three existing datasets of text-molecule pair: PubchemSTM curated by [30], ChBEI-20 [9], and PCdes [55]. Only a train split for each dataset was used for the training, and pairs that appear in the test set for the experiments in the Results are additionally removed. We also used 100,000 molecules from PubChem without any text descriptions, which helps the model learn the common distribution of molecules. When these unlabeled data were fed into the training model, we used a pre-defined null text which represents the absence of text condition.
The latent diffusion model was trained with the training loss suggested in ADM [8]. To take advantage of classifier-free guidance, we randomly replaced 3% of the given text condition with a null text during the training. More detailed training hyperparameters can be found in Appendix A. Unless specifically mentioned, the sampling iteration in the inference stage used DDIM-based [45] 100 sampling steps with a classifier-free guidance scale of 5.0.
@@ 4.1 Text-conditioned molecule generation
In this section, we evaluated the trained LDMol’s ability to generate molecules that agree with the given natural language conditions. We designed two scenarios: a benchmark study to see if the model can find the output that satisfies very long and specific conditions, and zero-shot text-to-molecule generation with unseen text conditions to see LDMol’s generalization ability and the learned alignment between text and molecules.
Quantitative benchmark study. We generated molecules with LDMol using the text captions in the ChEBI-20 test set and compared them with the ground truth. Here, we assume that the given description is long and detailed so that the ideal generative model can figure out the ground truth. On top of the training with molecule-text pairs from various sources, we additionally fine-tuned the model with the ChEBI-20 training set and 10,000 unlabeled molecules. The metrics we’ve used are as follows: SMILES validity, BLEU score [34] and Levenshtein distance between two SMILES, Tanimoto similarity [2] between two SMILES with three different fingerprints (MACCS, RDK, Morgan), the exact match ratio, and Frechet ChemNet Distance (FCD) [37].
We note that since our diffusion model sampling starts from randomly sampled noise from a prior distribution, the output from deterministic DDIM sampling still has stochasticity. However, the performance difference was marginal between starting from random noise and starting from a fixed point (e.g., zero vector). Since the given condition is more specific and detailed than zero-shot case studies, we used the lower classifier-free guidance scale of 3.5.
Table 1: Benchmark results of text-to-molecule generation on ChEBI-20 test set. The best performance for each metric was written in bold, and the second-best performance was underlined. The results for baseline models are taken from  [13].
@@
Transformer 0.906 0.499 57.660 0.480 0.320 0.217 0.000 11.32
SciBERT [3] 0.915 0.459 24.458 0.721 0.588 0.529 0.005 -
T5base [38] 0.660 0.765 24.950 0.731 0.605 0.545 0.069 2.48
MolT5base [10] 0.772 0.769 24.458 0.721 0.588 0.529 0.081 2.18
GIT-Mol [29] 0.928 0.756 26.315 0.738 0.582 0.519 0.051 -
TGM-DLM [13] 0.871 0.826 17.003 0.854 0.739 0.688 0.242 0.77
LDMol 0.928 0.900 9.259 0.931 0.860 0.820 0.363 0.25
Table 1 contains the performance of LDMol and other baselines for text-to-molecule generation on the ChEBI-20 test set. Including both transformer-based autoregressive models and diffusion-based models, LDMol outperformed the existing models in almost every metric. While one model showed similar validity to ours, it had a poor agreement with the ground truth compared to LDMol, with an exact match of 0.051. One thing to emphasize is that MolT5 uses the same text encoder with LDMol, yet there’s a significant performance difference between the two models. We believe this is because our constructed latent space is much easier to align with the same textual information, compared to the token sequence used by transformer-based models.
Figure 3: Examples of the generated molecules by LDMol with various text conditions. Validity on 1,000 generated samples is included in the lower left corner of each box.
Zero-shot case studies. To demonstrate the generalization ability of LDMol, we analyzed the generated output with several hand-written text conditions. These input prompts were not contained in the training data and were relatively vague and high-level so that many different molecules could satisfy the condition.
Figure 3 shows the samples of generated molecules from LDMol with several input prompt examples. We found that LMDol can generate molecules with high validity that follow the various levels of input conditions for specific atoms(a), functional groups(b), molecular substructure(c), compound class(d), and substance names(e). The validity was calculated by the number of valid SMILES over the 1,000 generated samples, which was above 0.9 for all five scenarios we’ve shown. Considering that these short, broad, and hand-written text conditions are distinct from the text conditions in the training dataset, we’ve concluded that our model is able to learn the general relation between natural language and molecules.
@@ 4.2 Applications toward downstream tasks
Well-trained diffusion models learned the score function of a data distribution, which enables high applicability to various downstream tasks. The state-of-the-art image diffusion models have shown their versatility in image editing [32, 16], classification [25], retrieval [21], inverse problems like inpainting and deblurring [7], image personalization [42], etc. To demonstrate LMDol’s potential versatility as a diffusion model, we applied the pre-trained LDMol to the molecule-to-text retrieval and text-driven molecule editing task.
Molecule-to-text retrieval. Our approach to molecule-to-text retrieval is similar to the idea of using a pre-trained diffusion model as a classifier [25]; LDMol takes each candidate text with a query molecule’s noised latent, and retrieves the text that minimizes the difference between output noise $\hat{\epsilon_{\theta}}$ and the injected noise $\epsilon$ (see Appendix B for more details). Since this process has randomness due to the stochasticity of $t$ and $\epsilon$, we repeated the same process $n$ times with resampled $t$ and $\epsilon$ and used a mean loss to minimize the performance variance.
Table 2: 64-way accuracy of molecule-to-text retrieval task. The custom test set and the baseline model results are taken from  [49]. For LDMol, $n$ is a number of iterations where $||\hat{\epsilon_{\theta}}-\epsilon||^{2}_{2}$ was calculated. The best performance for each task is written in bold.
Model PCdes test set custom test set from Su et al.
@@
SciBERT [3] 50.38 82.57 1.38 1.38
KV-PLM [55] 55.92 77.92 1.37 1.51
MoMu-S [49] 58.64 80.62 39.50 45.66
MoMu-K [49] 58.74 81.09 39.07 46.21
LDMol($n$=10) 52.28 89.36 57.90 78.16
LDMol($n$=25) 55.94 90.82 63.27 81.13
We measured a 64-way retrieval accuracy of LDMol using two different test sets: PCdes test split and another text-molecule retrieval dataset separately curated by  [49]. Table 2 contains the text-to-molecule accuracy of LDMol with other baseline models. Only one randomly selected sentence in each candidate description was used for the retrieval in the “sentence-level" column, and the entire descriptions were used for the “paragraph-level" column. LDMol achieved a higher performance in three out of four scenarios compared to the previously presented models and became more accurate as the number of function evaluations increased. It maintained its performance on a relatively out-of-distribution test set with minimal accuracy drop. The actual examples from the retrieval result can be found in Appendix C.
Text-driven molecule editing. We applied a method of Delta Denoising Score (DDS) [16], which was originally suggested for text-guided image editing, to see whether LDMol can be used to optimize a source molecule to match a given text. Using two texts that describe the source data $z_{src}$ and the desired target, DDS presents how a text-conditioned diffusion model can modify $z_{src}$ to a new data $z_{tgt}$ that follows the target text description (more detailed procedures can be found in Appendix B).
Figure 4: Examples of text-guided image editing with LDMol. The difference between the source text and the target text, and the corresponding region is colored in purple.
We imported a method of DDS on LDMol’s molecule latent to edit a given molecule to match the target text, where the case studies are shown in Figure 4., The editing results successfully modified the input molecule towards the target text with minimal corruption of the unrelated region. We emphasize that this kind of direct data optimization is possible because of the continuous and non-restricted latent domain from the LDMol’s encoder.
@@ 4.3 Ablation studies
We’ve conducted ablation studies for each design choice of the proposed LDMol to analyze and emphasize their role. We measured the SMILES validity and the exact match ratio of text-to-molecule generation on the ChEBI-20 test set, with three models pre-trained with the same number of DiT training iterations.
Table 3: Quantitative results of the ablation study.
@@
LDMol 0.922 0.303
w/o contrastive encoder pre-training $0$
w/o compression layer 0.110 0.003
The result of the ablation study is listed in Table 3. When the molecule latent space was constructed by a naive autoencoder (i.e. the SMILES encoder is not pre-trained with the proposed method), the diffusion model completely failed to learn the data distribution to generate valid SMILES. A similar problem occurred when we didn’t introduce a compression layer to reduce the dimension of the latent space, meaning that the diffusion model could not properly model the data distribution in the given latent space.
@@ 5 Conclusion
In this work, we presented a text-conditioned molecule diffusion model LDMol that runs on a chemical latent space reflecting structural information. By introducing the deeply studied paradigm of the latent diffusion model in general domains into the molecule domain with minimal modifications, LDMol retains many advanced attributes of diffusion models that enable various applications.
Despite the noticeable performances of LDMol, it still has limitations that can be improved; Since the text captions of our obtained paired data mostly describe structural features, LDMol still struggles to follow some text conditions such as biological properties. The current suggested text-guided molecule editing performance is relatively sensitive to hyperparameters as the imported method assumes broader and more substantial knowledge that emerged from the data generation. Nonetheless, we expect that the LDMol’s performance could be improved further with the emergence of richer text-molecule pair data and more powerful text encoders, with possible future works for the construction of a better latent space.
We believe that our approach could inspire tackling various chemical generation tasks using latent space, not only text-conditioned but also many more desired properties such as biochemical activity. Especially, we expect LDMol could be a starting point to bridge achievements in the state-of-the-art diffusion model into the chemical domain.
Appendix A Model hyperparameters and training setup
Table 4: Our choice of the model hyperparameters and training setup.

### v2

@@ 3.2 Text-conditioned latent diffusion model
As shown in Figure 2-(c), our diffusion model learns the conditional distribution of the SMILES latent $z$ whose dimension is $[L\times d_{z}]$. In the training phase, a molecule in the training data is mapped to the latent $z$ and applied a forward noising process into $z_{t}$ with randomly sampled diffusion timestep $t$ and injected noise $\epsilon$. A diffusion model predicts the injected noise from $z_{t}$, conditioned by the paired text description via a frozen external text encoder. In the inference phase, the diffusion model iteratively generates a new latent sample $z^{\prime}$ from a given text condition, which is then decoded to a molecule via the SMILES decoder.
Since most contributions to diffusion models were made in the image domain, most off-the-shelf diffusion models have the architecture of convolution-based Unet (Ronneberger et al., 2015). However, introducing the spatial inductive bias of Unet cannot be justified for the latent space of our encoder. Therefore we employed DiT (Peebles & Xie, 2023) architecture, one of the most successful approaches to transformer-based diffusion models for more general data domain. Specifically, we utilized a DiTbase model with minimal modifications to handle text conditions with cross-attention, where more details can be found in Section A.1.
Text-based molecule generation requires a text encoder to process natural language conditions. Existing text-based molecule generation models trained their text encoder from scratch (Pei et al., 2023), or utilize a separate encoder model pre-trained on scientific domain corpora (Beltagy et al., 2019). In this work, we took the encoder part of MolT5large (Edwards et al., 2022) as our text encoder.
@@ 3.3 Implementation details
The pre-training of the SMILES encoder and the corresponding decoder was done with 10,000,000 general molecules from PubChem (Kim et al., 2023). To ensure enough batch size for negative samples (He et al., 2020), we build a memory queue that stores $Q$ recent input and use them for the encoder training. The SMILES tokenizer vocabulary consists of 300 tokens, which were obtained from the pre-training data SMILES corpus using the BPE algorithm (Gage, 1994). We only used SMILES data that does not exceed a fixed maximum token length $L$.
To train the text-conditioned latent diffusion model, we gathered three existing datasets of text-molecule pair: PubchemSTM curated by Liu et al. (2023a), ChEBI-20 (Edwards et al., 2021), and PCdes (Zeng et al., 2022). Only a train split for each dataset was used for the training, and pairs that appear in the test set for the experiments are additionally removed. We also used 10,000 molecules from ZINC15 (Sterling & Irwin, 2015) without any text descriptions, which helps the model learn the common distribution of molecules. When these unlabeled data were fed into the training model, we used a pre-defined null text which represents the absence of text condition.
The latent diffusion model was trained with the training loss suggested by Dhariwal & Nichol (2021). To take advantage of classifier-free guidance (Ho & Salimans, 2021), we randomly replaced 3% of the given text condition with the null text during the training. The sampling iteration in the inference stage used DDIM-based (Song et al., 2021a) 100 sampling steps with a classifier-free guidance. More detailed training hyperparameters can be found in Appendix A.2.
@@ 4.1 Text-conditioned molecule generation
In this section, we evaluated the trained LDMol’s ability to generate molecules that agree with the given natural language conditions. First, we generated molecules with LDMol using the text captions in the ChEBI-20 test set and compared them with the ground truth. The metrics we’ve used are as follows: SMILES validity, BLEU score (Papineni et al., 2002) and Levenshtein distance between two SMILES, Tanimoto similarity (Bajusz et al., 2015) between two SMILES with three different fingerprints (MACCS, RDK, Morgan), the exact match ratio, and Frechet ChemNet Distance (FCD) (Preuer et al., 2018).
We note that since our diffusion model sampling starts from randomly sampled noise from a prior distribution, the output from deterministic DDIM sampling still has stochasticity. However, with a highly detailed text prompt that can specify a single correct answer, we discovered that LDMol was powerful enough to outperform many autoregressive models with randomly sampled initial points. Also, the performance difference was marginal between starting from random noise and starting from a fixed point (e.g., zero vector). We tested different scales for the classifier-free guidance scale $\omega$ in the sampling process and found $\omega=2.5$ works best (See Section B.1).
Table 1: Benchmark results of text-to-molecule generation on ChEBI-20 test set. The best performance for each metric was written in bold. The “Family” column denotes whether the model is AR(autoregressive model) or DM(diffusion model).
@@
Transformer AR 0.906 0.499 57.660 0.480 0.320 0.217 0.000 11.32
GIT-Mol (Liu et al., 2024a) AR 0.928 0.756 26.315 0.738 0.582 0.519 0.051 -
T5base (Raffel et al., 2020) AR 0.660 0.765 24.950 0.731 0.605 0.545 0.069 2.48
MolT5base (Edwards et al., 2022) AR 0.772 0.769 24.458 0.721 0.588 0.529 0.081 2.18
T5large AR 0.902 0.854 16.721 0.823 0.731 0.670 0.279 1.22
MolT5large AR 0.905 0.854 16.071 0.834 0.746 0.684 0.311 1.20
MolXPT (Liu et al., 2023b) AR 0.983 - - 0.859 0.757 0.667 0.215 0.45
bioT5 (Pei et al., 2023) AR 1.000 0.867 15.097 0.886 0.801 0.734 0.413 0.43
bioT5+ (Pei et al., 2024) AR 1.000 0.872 12.776 0.907 0.835 0.779 0.522 0.35
TGM-DLM (Gong et al., 2024) DM 0.871 0.826 17.003 0.854 0.739 0.688 0.242 0.77
LDMol DM 0.941 0.926 6.750 0.973 0.950 0.931 0.530 0.20
Table 1 contains the performance of LDMol and other baselines for text-to-molecule generation on the ChEBI-20 test set. Including both transformer-based autoregressive models and diffusion-based models, LDMol outperformed the existing models in almost every metric. While few models showed higher validity than ours, they showed a lower agreement between the output and the ground truth, which we insist is a more important role of generative text-to-molecule models. One thing to emphasize is that MolT5large uses the same text encoder with LDMol, yet there’s a significant performance difference between the two models. We believe this is because our constructed latent space is much easier to align with the same textual information, compared to the token sequence used by transformer-based models.
Figure 4: Examples of the generated molecules by LDMol with various text conditions, with validity on 1,000 generated samples.
To demonstrate the LDMol’s molecule generalization ability with more broad and general text inputs, we analyzed the generated output with several hand-written prompts. These input prompts were not contained in the training data and were relatively vague and high-level so that many different molecules could satisfy the condition. Figure 4 shows the samples of generated molecules from LDMol with several input prompt examples. We found that LMDol can generate molecules with high validity that follow the various levels of input conditions for specific atoms(a), functional groups(b), molecular substructure(c), compound class(d), and substance names(e). The validity was calculated by the number of valid SMILES over 1,000 generated samples, above 0.9 for most scenarios we tested. Considering that these short, broad, and hand-written text conditions are distinct from the text conditions in the training dataset, we’ve concluded that our model is able to learn the general relation between natural language and molecules.
@@ 4.2 Applications toward downstream tasks
Well-trained diffusion models learned the score function of a data distribution, which enables high applicability to various downstream tasks. The state-of-the-art image diffusion models have shown their versatility in image editing (Meng et al., 2022; Hertz et al., 2023), classification (Li et al., 2023), retrieval (Jin et al., 2023), inverse problems like inpainting and deblurring (Chung et al., 2023), image personalization (Ruiz et al., 2023), etc. To demonstrate LMDol’s potential versatility as a diffusion model, we applied the pre-trained LDMol to the molecule-to-text retrieval and text-guided molecule editing task.
Molecule-to-text retrieval. Our approach to molecule-to-text retrieval is similar to the idea of using a pre-trained diffusion model as a classifier (Li et al., 2023): LDMol takes each candidate text with a query molecule’s noised latent, and retrieves the text that minimizes the noise estimation error $||\hat{\epsilon}_{\theta}-\epsilon||^{2}_{2}$ between the injected noise $\epsilon$ and the predicted noise $\hat{\epsilon}_{\theta}$(see Appendix A.3 for more details). Since this process has randomness due to the stochasticity of $t$ and $\epsilon$, we repeated the same process $n$ times with resampled $t$ and $\epsilon$ and used a mean error to minimize the performance variance.
Table 2: 64-way accuracy in % on molecule-to-text retrieval task. For LDMol, $n$ is a number of iterations where $||\hat{\epsilon}_{\theta}-\epsilon||^{2}_{2}$ was calculated. The best performance for each task is written in bold.
Model PCdes test set MoMu test set
@@
SciBERT (Beltagy et al., 2019) 50.4 82.6 1.38 1.38
KV-PLM (Zeng et al., 2022) 55.9 77.9 1.37 1.51
MoMu-S (Su et al., 2022) 58.6 80.6 39.5 45.7
MoMu-K 58.7 81.1 39.1 46.2
MoleculeSTM (Liu et al., 2023a) - 81.4 - 67.6
MolCA (Liu et al., 2024b) - 86.4 - 73.4
LDMol($n$=10) 60.7 90.2 66.4 84.8
LDMol($n$=25) 62.2 90.3 78.4 87.1
We measured a 64-way in-batch retrieval accuracy of LDMol using two different test sets: PCdes test split and MoMu retrieval dataset curated by Su et al. (2022), where the result with other baseline models are listed in Table 2. Only one randomly selected sentence in each candidate description was used for the retrieval in the “sentence-level” column, and all descriptions were used for the “paragraph-level” column. LDMol achieved a higher performance in all four scenarios compared to the previously presented models and maintained its performance on a relatively out-of-distribution MoMu test set with minimal accuracy drop. LDMol became more accurate as the number of function evaluations increased, and the improvement was more significant at the sentence-level retrieval and out-of-distribution dataset. The actual examples from the retrieval result can be found in Appendix B.2.
Figure 5: (a) Hit ratio of molecule editing by LDMol and MoleculeSTM (Liu et al., 2023a) in eight scenarios. Following MoleculeSTM, each scenario was applied to 200 randomly sampled molecules from ZINC15, and the mean and standard variation on three separate runs were plotted. In the figure, we omitted “This molecule” in front of the actual prompts, and abbreviated “hydrogen bonding” to “H-bond”. (b) Examples of text-guided molecule editing with LDMol. The difference between the source text and the target text, and the corresponding region is colored in purple.
Text-guided molecule editing. We applied a method of Delta Denoising Score (DDS) (Hertz et al., 2023), which was originally suggested for text-guided image editing, to see whether LDMol can be used to optimize a source molecule to match a given text. Using two text prompts that describe the source data $z_{src}$ and the desired target, DDS presents how a text-conditioned diffusion model can modify $z_{src}$ into a new data $z_{tgt}$ that follows the target text prompt (more detailed procedures can be found in Appendix A.3).
We imported a method of DDS on LDMol’s molecule latent to edit a given molecule to match the target text, with several prepared editing prompts that require the model to change certain atoms, substructures, and intrinsic properties from the source molecule. Figure 5-(a) shows that LDMol had comparable performance with a previously suggested text-guided molecule editing model (Liu et al., 2023a), with a higher hit ratio in five out of eight scenarios. Actual case studies with editing sources and results are listed in Figure 5-(b), where the editing results successfully modified the input molecule towards the target prompt with minimal corruption of the unrelated region.
@@ 4.3 Ablation studies
We’ve conducted ablation studies for each design choice of the proposed LDMol to analyze and emphasize their role. Each model is pre-trained with the same number of DiT training iterations. The results of the ablation studies are listed in Table 3.
Table 3: Quantitative results of the ablation studies.
models Autoencoder Text-to-molecule generation
@@
LDMol w/o contrastive learning(i.e. naive AE) 1.000 0.019 0.000 58.60
LDMol w/o compression layer 0.964 0.022 0.000 67.93
LDMol w/o stereoisomer hard-negative 0.891 0.939 0.278 0.24
naive AE + KL regularization($\beta=0.001$) 0.999 0.847 0.492 0.34
LDMol 0.983 0.941 0.530 0.20
When we remove the contrastive encoder pre-training objective and construct the molecule latent space with a naive autoencoder, the diffusion model completely fails to learn the latent distribution to generate valid SMILES. A similar problem occurred when we didn’t introduce a compression layer since the dimension of the latent space was too big for the diffusion model to learn. When stereoisomers were not utilized as hard negative samples in the contrastive encoder training, the constructed latent space was not detailed enough to specify the input, which degraded the reconstruction accuracy of the autoencoder.
An interesting case is the comparison against the autoencoder with KL-divergence loss: it has reconstructible latent space and showed a text-to-molecule generation match ratio of 0.492, which outperforms several baseline models in Table 1. This performance shows the ability of diffusion models on the continuous data domain which can be comparable to the autoregressive models. However, its overall metric is still lower than the proposed LDMol, with especially low validity and FCD. We insist that this gap comes from the structurally informative latent space of LDMol which is easier for the diffusion model to learn the correlation between the latent space and the condition.
@@ 5 Conclusion
In this work, we presented a text-conditioned molecule diffusion model LDMol that runs on a chemical latent space reflecting structural information. By introducing the deeply studied paradigm of the latent diffusion model into the molecule domain with minimal modifications, LDMol retains many advanced attributes of diffusion models that enable various applications.
Despite the noticeable performances of LDMol, it still has limitations that can be improved: Since the text captions of our obtained paired data mostly describe structural features, LDMol still struggles to follow some text conditions such as complex biological properties. Nonetheless, we expect that the LDMol’s performance could be improved further with the emergence of richer text-molecule pair data and more powerful text encoders. Moreover, combining physiochemical and biological annotations on top of the structurally informative latent space is a promising future work that can ease the connection between molecules and text conditions.
We believe that our approach could inspire tackling various chemical generation tasks using latent space, not only text-conditioned but also many more desired properties such as biochemical activity. Especially, we expect LDMol to be a starting point to bridge achievements in the state-of-the-art diffusion model into the chemical domain.
Appendix A Experimental details
A.1 DiT block for SMILES latent diffusion model with text conditions
The DiT block architecture of the class-conditioned image diffusion model published by Peebles & Xie (2023) is shown in Figure 6-(a). The noised input image latent $z_{t}$ is passed through a patch embedding layer and spatially flattened to be fed into the DiT block. The condition embedding $y$ and diffusion timestep embedding $t$ are incorporated into the model prediction via adaptive layer norm. The dimension of $t$ and $y$ are both $[B\times F]$, where $B$ is the batch size and $F$ is the number of features.
In the case of LDMol, the input latent $z_{t}$ with dimension $[B\times L\times F]$ is already spatially one-dimensional, we simply pass it through a linear layer to prepare DiT block input. Also, the text condition feature we’ve used has a much higher dimension of $[B\times L^{\prime}\times F]$ where $L^{\prime}$ is the token length of the text condition. Therefore, we stacked a cross-attention layer for text condition features after each self-attention layer, as shown in Figure 6-(b).
Figure 6: Input embedding layer and DiT blocks in the (a) originally published DiT and (b) LDMol.
A.2 Model hyperparameters and training setup
Detailed hyperparameters on the model architecture are listed in Table 4, with settings on the model training procedure.
Table 4: The choice of the model hyperparameters and training setup.